5. Calculate the percent copper and percent chlorine for the following compound:

Cu2Cl

Group of answer choices

50.00% Cu, 50.00% Cl

28.55% Cu, 71.45% Cl

71.45% Cu, 28.55% Cu

78.19% Cu, 21.81% Cl

21.81% Cu, 78.19% Cl

6. Calculate the percent copper and percent chlorine for the following compound:

CuCl2

Group of answer choices

78.19% Cu, 21.81% Cl

21.81% Cu, 78.19% Cl

47.26% Cu, 52.74% Cl

52.74% Cu, 47.26% Cl

25% Cu, 75% Cl

7. Using the following data and the percents calculated in the previous questions #5 & #6, determine the empirical formula for copper chloride.

mass of evaporating dish and copper = 25.097 g

mass of evaporating dish = 23.442 g

mass of copper chloride = 3.502 g

Group of answer choices

Cu2Cl

CuCl2

8.

In this experiment you placed zinc into a copper chloride solution. What would you classify this reaction?

Group of answer choices

single replacement

double replacement

decomposition

combination

combustion

9. Between the copper chloride and zinc as the reactants of this experiment. Which of these two reactants was present in excess?

Group of answer choices

copper chloride

zinc
Answer #1

5. Cu2Cl

Mass of Compound = 2 * 63.5 + 35.5 = 162.5 amu

Mass of Cu = 2 * 63.5 = 127

% mass Cu = 127 / 162.5 * 100 = 78.19%

% Cl = 100 - 78.19 = 21.1%

6.

CuCl2

Mass of Compound = 63.5 + 2 * 35.5 = 134.5 amu

Mass of Cu = 63.5

% mass Cu = 63.5 / 134.5 * 100 = 47.26%

% Cl = 100 - 47.26 = 52.74%

8. Single replacement

9. Copper chloride

7. depends on your experiment